OpenStudy (anonymous):

square root property and completing the square:- x^2+3x-4=0

OpenStudy (anonymous):

(x+4)(x-1)

OpenStudy (anonymous):

\[x^2+3x-4=0\] \[(x+4)(x-1)=0\] Therefore x = -4 or x =1. (That's by factoring.) The square root property basically says\[\sqrt{a^2}=\pm{a}\] To complete the square, you want all non-x coefficients on one side and all x-coeffs. on the other.\[x^2+3x=4\]Then you add (b/2)^2 to both sides, using\[ax^2+bx+c=0\]as a reference, so you get\[(3^2/4)=9/4\]\[x^2+3x+9/4=4+9/4\]And now it is possible to factor the equation so it looks like\[(x+a)^2\]